There is a point charge Q1 at x= 0 and a charge Q2 at x=d. What is the relationship between these charges if the resulting electric field is zero at the following points: a) x= d/2 b) x= 2d c) x= -d/2?

1
Expert's answer
2022-02-21T14:56:49-0500

"E_i=\frac{kq_i}{r_i^2},"

a)

"\frac{q_2}{q_1}=1,"

b)

"\frac{q_2}{q_1}=-0.25,"

c)

"\frac{q_2}{q_1}=-9."